OpenStudy (anonymous):

Find the greatest common factor of these two expressions. 8u^6v^8 and 28u^7v^2x^3

OpenStudy (mathstudent55):

First, find the GCF of the coefficients (the numbers) 8 and 28.

OpenStudy (mathstudent55):

\(8u^6v^8\) \(28u^7v^2x^3\) Show every factor of the two terms: \(8u^6v^8 = 2 \times 2 \times 2 \times uuuuuu \times vvvvvvvv\) \(28u^7v^2x^3 = 2 \times 2 \times 7 \times uuuuuuu \times vv \times xxx\)

OpenStudy (anonymous):

?

OpenStudy (usukidoll):

mathstudent55 expanded everything out for you... you just have to figure out what the two of those expressions have in common

OpenStudy (mathstudent55):

The GCF is the product of all factors both terms have in common: \(8u^6v^8 = \color{red}{2 \times 2} \times 2 \times \color{red}{uuuuuu} \times \color{red}{vv}vvvvvv\) \(28u^7v^2x^3 = \color{red}{2 \times 2} \times 7 \times \color{red}{uuuuuu}u \times \color{red}{vv} \times xxx\)

OpenStudy (mathstudent55):

You see above the common factors in red.